Really need help with this problem please
Jim wants to plan a meal with 134 grams of carbohydrates and 1180 calories if green beans have 7 grams of carbohydrates and 30 calories per half-cup serving and fried shrimp have 9 grams of carbohydrates and 190 calories per three-ounce serving, how many servings of green beans and shrimp should he use?
A) 9 half cup servings of green beans and 7 there ounce servings of shrimp
B)4 half cup servings of green beans and 14 there ounce servings of shrimp
C)7 half cup servings of green beans and 9 there ounce servings of shrimp
D)14 half cup servings of green beans and 4 there ounce servings of shrimp

Found 2 solutions by Boreal, macston:
Answer by Boreal(15235)

The easiest way to do this is to check each answer.
A has 63/270+9/190=72/460
B has 28/120 + 126/1710=154/1830
C has 49/210 + 81/1710 too much
D has 98/420 + 36/760=134/1180. D works.
7x+9y=134
30x+190y=1180
210x+270y=4020, multiplying by 30
-210x-1330y=-8260, multiplying by minus 7
-1060=-4240
y=4
4 three ounce servings of shrimp, D.







Answer by macston(5194)

Easiest to check is carbs:
A) 63g+63g=126g
B) 28g+128g=156g
C) 49g+81g=130g
D) 98g+36g=134g
Choice D) is the only one to meet carbohydrate requirement, so we check the calories: 420 cal + 760 cal=1180 calories, so calories also meet requirements.
ANSWER: The correct choice is (D).